OpenStudy (anonymous):

Jodie has a rectangular room with dimensions 6 feet 8 inches × 12 feet. Jodie decided to draw a rectangular sketch of her room with dimensions 5 inches × 9 inches. Is her sketch similar to her actual room?

OpenStudy (anonymous):

1feet = 12inches 6feet=12*6=72inches 6feet 8inches=72+8=80inches now for the ordinate, 12feet=12*12=144inches now to check if 5in*9in is in proportion with 80in*144in 80/5 should be equal to 144/9 which is not the case so they are not in proporition anf this her sketch is not similar to her actual room

OpenStudy (anonymous):

it yes

OpenStudy (anonymous):

80-5=75 and 144-9=135 75 is divisible by 5 and 135 is divisible by 9..so the ans is yes

OpenStudy (anonymous):

sorry i cant xplain coz m confused myself.. lol..XD

OpenStudy (anonymous):

Following from deeprony7: 80/5 = 16 144/9=16 hence 80/5=144/9 yes it's proportional which means it is similar =D

OpenStudy (anonymous):

ooops.. need to redo my kindergarden..lol

OpenStudy (anonymous):

The shadow of a 5-foot tall post is 20 feet long. At the same time a tree casts a shadow that is 56 feet long. What is the height of the tree?